Aqueous solution of HCL is diluted by five times. What will be the change in pH?

Grade:12

1 Answers

Sunil Kumar FP
9 years ago
ph is measured in terms of the activity of the H+ io.
pH=-log(H+)
on dilutimg the solution 5 times means increasing the volume from v to 5v
let the number of mole of H+ ion be x
initial pH at a volume V=-log[H+] =-log[x/v] =pH1
now on dilution new pH= -log[H+]=-log [x/5v]=pH2
change in ph =pH2 -pH1 =-log(x/5v)-(-log(x/v)) =log(x/v)-log(x/5v) =log 5 =.6989